OpenStudy (anonymous):

Variables on both sides. Solve each equation. -5v-4v=-6v-6v-12

Nnesha (nnesha):

combine like terms both side which is -5v and -4v on left side and -6v and -6v on right side

OpenStudy (anonymous):

Questions like this involve recognising "like terms". If two or more numbers have the same variable "attached" to them (multiplied with them) it means you can add and subtract those numbers! For example, -6v - 6v is just like doing -6-6, but you just let the v join along. So something like 2x + 3x = 5x, just like 2+3=5. In your situation, you need to: 1. "Collect" like terms. 2. Move all the variables to one side of the equation (this can be achieved by adding or subtracting the same number from both sides of the equation). 3. Dividing both sides of the equation by whatever number is "attached" to the variable, so that that number disappears and you have the variable by itself.

OpenStudy (anonymous):

For example, if you had 7x - 2x = 3x + 4 Your next step would be to collect like terms on both sides, and get 5x = 3x + 4 Then you'd need to make it so your variable is only on ONE side of the equation - let's make it the left hand side. That means we need to take away 3x from both sides. 5x - 3x = 3x + 4 - 3x (we're doing the same thing to both sides of the equation here!) 2x = 4 Now we divide both sides by 2 to get x = 2. Try a similar approach for your situation.